Some students found a pirate treasure map with the following clues: "Take ye 20 paces East of the old oak tree; then 15 North and 18 West. Walk 9 paces North and another 5 East and here ye find me tresure." How many paces, in a straight line, was the treasure from the oak tree?
Answer by stanbon(75887)

Some students found a pirate treasure map with the following clues:
Let the oak tree be at (0,0)
------------------
"Take ye 20 paces East of the old oak tree
You are now at (20,0)
------------------
; then 15 North and 18 West
you are now at (2,0)
--------------------
. Walk 9 paces North
You are now at (2,9)
--------------------
and another 5 East
You are now at (7,9)
--------------------
and here ye find me tresure."
------------------------------
How many paces, in a straight line, was the treasure from the oak tree?
Find the distance from (0,0) to (7,9)
d = sqrt(7^2+9^2) = sqrt(130) ft.
=======================================
Cheers,
Stan H.
=============